OpenStudy (anonymous):

Let A be the area of a circle with radius r. If dr/dt=3, find dA/dt when r=3.

OpenStudy (anonymous):

Since A = pi(r^2), taking the derivative of both sides gives you dA/dt = pi (2r)(dr/dt). You have to use the chain rule when taking the derivative of r. Now just plug in your values. dr/dt = 3, r = 3, so now you have dA/dt = pi (2(3))(3)=18pi

OpenStudy (anonymous):

thanks for the step by step

OpenStudy (anonymous):

another problem i had trouble with was Let y=3x^2+6x+3 Find the differential dy when x=2 and dx =0.1 Find the differential dy when x=2 and dx =0.2 for the derivative of y = 6x+6 \[dy=(6x+6)dx\] from there on i plugged in 6(2)+6(0.1) i got 12.6 but online says the answer is wrong

OpenStudy (anonymous):

ah, you were close, you did the Calculus part right (the derivative), but you made a small mistake while calculating. it would be dy =(6x+6)dx = (6(2)+6)(0.1) = (18)(0.1) = 1.8

OpenStudy (anonymous):

ohhh okay i see so the dx is multiplied separately!! thanks i was wondering what was wrong